OpenStudy (uscrnamc):

let f(x) = 8^x and g(x) = 8(^x+5) + 1 which transformations are need to transform the graph of f(x) to the graph g(x)?

OpenStudy (uscrnamc):

answers

OpenStudy (jdoe0001):

\(\bf \qquad \qquad \qquad \qquad \textit{function transformations} \\ \quad \\ \begin{array}{rllll} % left side templates f(x)=&{\color{purple}{ A}}({\color{blue}{ B}}x+{\color{red}{ C}})+{\color{green}{ D}} \\ \quad \\ y=&{\color{purple}{ A}}({\color{blue}{ B}}x+{\color{red}{ C}})+{\color{green}{ D}} \\ \quad \\ f(x)=&{\color{purple}{ A}}\sqrt{{\color{blue}{ B}}x+{\color{red}{ C}}}+{\color{green}{ D}} \\ \quad \\ f(x)=&{\color{purple}{ A}}\mathbb{R}^{{\color{blue}{ B}}x+{\color{red}{ C}}}+{\color{green}{ D}} \end{array}\qquad \begin{array}{llll} % right side info \bullet \textit{ stretches or shrinks horizontally by } {\color{purple}{ A}}\cdot {\color{blue}{ B}}\\ \bullet \textit{ horizontal shift by }\frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\\ \qquad if\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\textit{ is negative, to the right}\\ \qquad if\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\textit{ is positive, to the left}\\ \bullet \textit{ vertical shift by }{\color{green}{ D}}\\ \qquad if\ {\color{green}{ D}}\textit{ is negative, downwards}\\ \qquad if\ {\color{green}{ D}}\textit{ is positive, upwards} \end{array} \\ \quad \\ -----------------------------\\ g(x)=8^{x+{\color{red}{ 5}}}+{\color{green}{ 1}}\) you tell us

OpenStudy (jdoe0001):

\(\large \bf g(x)=8^{{\color{blue}{ 1}} x+{\color{red}{ 5}}}+{\color{green}{ 1}}\) just clarifying a bit more, in case needed

OpenStudy (3mar):

Just what @jdoe0001 has written,,,, What do you think, @uscrnamc? How many steps in the horizontal direction \(\color{red}x\) and how many in the vertical direction \(\color{red}y\)?

OpenStudy (uscrnamc):

the only one i really know that i did by myself is it went up 5 units

OpenStudy (3mar):

Just what @jdoe0001 has written,,,, What do you think, @uscrnamc? How many steps in the horizontal direction \(\color{red}x\) and how many in the vertical direction \(\color{red}y\)?

OpenStudy (jdoe0001):

the 5 is "grouped" with the "x", thus is a horizontal shift, doesn't "go up by 5 units" going up means, a vertical shift